### v3

# RADLADS: Rapid Attention Distillation to Linear Attention Decoders at Scale

Abstract
We present Rapid Attention Distillation to Linear Attention
Decoders at Scale (RADLADS), a protocol for rapidly converting softmax attention transformers into linear attention decoder models, along with two new RWKV-variant architectures, and models converted from popular Qwen2.5 open source models in 7B, 32B, and 72B sizes. Our conversion process requires only 350-700M tokens, less than 0.005% of the token count used to train the original teacher models. Converting to our 72B linear attention model costs less than $2,000 USD at today’s prices, yet quality at inference remains close to the original transformer. These models achieve state-of-the-art downstream performance across a set of standard benchmarks for linear attention models of their size. We release all our models on HuggingFace under the Apache 2.0 license, with the exception of our 72B models which are also governed by the Qwen License Agreement. 111Training Code: https://github.com/recursal/RADLADS-paper Models:
https://huggingface.co/collections/recursal/radlads-6818ee69e99e729ba8a87102
1 Introduction
Linear attention transformer variants employing compressive state have begun to match or even exceed traditional softmax attention based transformers in quality across many metrics. (Peng et al., 2025) This is important for long sequences at inference time, as linear attention is computable in O(1) time per token instead of O(N) time for softmax transformers, and avoids the expensive memory bandwidth usage of a Key Value cache. (Arora et al., 2023) However, the cost of training usable large models at scale is cost prohibitive for all but the largest organizations. This is especially true for large language models, where SoTA results have often required training on over ten trillion tokens of data. (Qwen et al., 2025) (Grattafiori et al., 2024)
Relative scorea Name Tokens Required Stages Lambada MMLU Others SUPRA (Mercat et al., 2024) 100B 1 0.913 0.216 0.869 LOLCats/Hedgehogb (Zhang et al., 2024a) 40M 2 - -0.288 0.635 MOHAWK (Bick et al., 2024b) 3-5B 3 0.938 -0.047 0.924 Mamba in the Llama (Wang et al., 2024) 20B 2 0.608 0.519 0.757 DiJiangb (Chen et al., 2024) 40B 1 - 0.720 - ARWKV (Yueyu et al., 2025) 60M/830M 2/3 0.893 0.801 0.935 Llamba (Bick et al., 2025) 8-12B 3 c0.951 c0.837 c0.987 RADLADS (ours) 350-700M 3 0.983 0.924 1.013
We present a distillation method that yields comparable performance to the original teacher model, while using only between 350-700 million tokens, less than 0.005% of the tokens required to pre-train SoTA transformers. This approach enables the creation of large linear attention language models without incurring the extreme costs of multi-trillion token training runs. It also opens up new avenues for researchers who work on the next generation of compressive state attention variants to test, train, and release models containing their new designs at scale.
While converting models, we found that pre-existing RWKV architectural choices were sometimes imperfectly matched to the needs of conversion. As a result, we developed two new architectures RAD-RWKV6 (”RADFinch”) and RAD-RWKV7 (”RADGoose”) based on RWKV-6 and RWKV-7 that allow conversion to proceed smoothly and the resultant model to be faster to inference than had we used the original RWKV-6 and RWKV-7 designs. We share these new architectural details, as well as some of the reasons behind the changes. We hope that other researchers may find this useful in converting transformers to their own linear attention designs.
Our main contributions are:
Details of the RADLADS conversion process, including stepwise instructions, specific hyperparameters, token counts, and dataset choice.
Two new simplified RWKV based architectures: RAD-RWKV6 (”RADFinch”) and RAD-RWKV7 (”RADGoose”), used to convert softmax attention efficiently to linear attention.
Public release of several converted Qwen models: QRWKV6-7B-Base, QRWKV6-7B-Instruct, QRWKV7-7B-Instruct, QRWKV6-32B-Instruct, QRWKV6-72B-Instruct, and QRWKV6-QwQ-32B.
Open source conversion code, allowing anyone to adapt the RADLADS process to a transformer of their choice and emit a highly performant converted model, including for other recurrent linear attention and/or compressive state architectures.
2 Background
There have been many attempts at efficiently converting pre-trained transformer models to pure linear attention or other recurrent compressive state architectures. Early work, like Gerstenberger et al. (2020), performs full model logit distillation to a freshly initialized student model, and requires a very long training cycle. T2R (Kasai et al., 2021) attempts to reduce this training time by keeping most of the original model intact. It swaps softmax attention for traditional linear attention with a learnable MLP feature map, and finetunes the resultant model on approximately 3.5 billion tokens of data. Other projects like DiJiang (Chen et al., 2024), XATL (Choi, 2024), and SUPRA (Mercat et al., 2024) combine these two techniques, but still require 100B tokens of training or more, and exhibit lacking performance on some popular benchmarks like MMLU.
More recently, new model architectures such as GSA (Zhang et al., 2024b) have been proposed, which are designed to adapt well to a SUPRA-like conversion process. Post-conversion downstream performance is improved, but these still require long training cycles.
Other works such as Mamba in the Llama (Wang et al., 2024) have focused on a pipeline of progressive distillation followed by supervised fine-tuning (Kim & Rush, 2016) and direct preference optimization (Rafailov et al., 2024). Mamba in the Llama uses 20B tokens of training. However, it focuses on hybrid models rather than pure RNNs, and performs poorly when the softmax attention is removed completely.
A recent development in the conversion process, featured in LOLCats (Zhang et al., 2024a) and MOHAWK (Bick et al., 2024a) has been to separate the conversion process into two phases: attention alignment and full model knowledge distillation. The first phase aligns attention scores and/or attention hidden states with the teacher model, and the second performs traditional logit distillation. MOHAWK still requires training on a large amount of data. LOLCats (Hedgehog version) requires only 40m tokens of training, but benchmark scores remain low, with MMLU accuracy lower than random guessing. To address this deficiency, they create a hybrid with full softmax attention in a sliding window (SWA), which improves scores but is no longer a purely recurrent model. We theorize that their low scores are due to the use of LoRAs for training, as well as vanilla linear attention rather than a more advanced recurrent architecture. This is most evident when examining their relative MMLU scores, which even with the inclusion of SWA remain low compared to ours.
We shared our early code, techniques, and 32B model with another team, who worked with those and subsequently published ARWKV (Yueyu et al., 2025), which converts Qwen2.5-7B-Instruct to use a standard RWKV-7 sequence mixer. As we discuss later, specific choices for weight transfer, hyperparameters, dataset, and architecture matter significantly. This is visible in the significantly higher downstream performance of our own Qwen2.5-7B-Instruct conversion, as shown in Tables 1 and 3.
3 Design Choices
Our work builds upon this varied foundation of historical attempts to convert softmax attention transformers to linear attention, and we attempt to use all of the best parts in a single process. We transfer attention weights, align attention hidden states, distill, finetune, develop a strong and appropriate RNN attention-replacement architecture, use a good dataset, and find the proper sequence of distillation hyperparameters. Combining all of these together results in lower training token counts while achieving a new state-of-the-art in pure RNN downstream performance. The reduction in token count allows us to affordably convert very large models, up to 72 billion parameters so far.
We find that choosing a good RNN architecture is an important part of any highly successful conversion process. T2R, SUPRA, and LOLCats all use traditional linear attention or make small changes (e.g. removing the denominator in favor of normalization). Other modern RNN architectures have been shown to be more expressive than linear attention, even when including such changes. (Sun et al., 2023; Yang et al., 2024; Peng et al., 2025) But choosing a strong architecture is not always enough. We initially used an unmodified RWKV-6 architecture in our experiments, and found disappointing performance. Removing the off-by-one decay and bonus by using a Gated Linear Attention kernel allowed the model to fit the original softmax attention hidden states much more closely during step 1. And we find that RWKV-7 fits even closer and more rapidly during this step, resulting in much lower distillation loss with even less compute.
Some components of modern RNN architectures that appear important for pretraining turned out to have little to no effect during conversion. And some have useful effects in one architecture but not others. One interesting example is RWKV tokenshift, a variety of 1D short convolution. It was useful in RAD-RWKV6, but conferred essentially no benefit in RAD-RWKV7. Another such case is gating, which was extremely beneficial only at full rank in RAD-RWKV6, but performed perfectly well in RAD-RWKV7 with reduced rank. Careful testing of each component can help discover the most efficient variant for a given architecture for use in softmax attention conversion.
Choice of dataset appears to matter quite a bit. After trying several custom datasets, as well as fineweb and fineweb-edu (Lozhkov et al., 2024), we settled on DCLM (Li et al., 2024b) for all our conversions. We theorize that the best choice of dataset may depend upon the teacher model’s pre-training data distribution. DCLM worked exceptionally well for us when converting Qwen models.
Training for the correct number of tokens and with good choices for learning rates was also very important, both for efficiency of training and for final downstream performance. More training was not always better. We arrived at our final hyperparameters after theorizing that most knowledge resides in the teacher model’s MLPs and embeddings. Therefore, we start at a high learning rate to rapidly align the attention hidden states, annealing to approximately the final learning rate that the teacher model saw during pretraining. We tried various learning schedules, and cosine worked best. We keep the learning rate fixed during steps 2 and 3. This is an attempt to avoid catastrophic forgetting or overly significantly changing the knowledge encoded in the original teacher MLPs.
4 Method
In RADLADS, our student model architecture is an exact copy of the teacher model architecture, with the exception of the attention blocks. These are replaced with a choice of purely recurrent attention-replacement time mixing blocks. In this paper we detail two such replacements, which are based on RWKV6-C2 and RWKV7, which we name RAD-RWKV6 (”RADFinch”) and RAD-RWKV7(”RADGoose”) respectively. Our student models retain the runtime repetition of keys and values from Grouped Query Attention (GQA) (Ainslie et al., 2023) in the teacher model, if present. Depending on the attention-replacement chosen, we optionally include rotary positional embeddings (RoPE) (Su et al., 2023) from the teacher model. In cases where we include RoPE, it is applied at exactly the same point as in the teacher model.
See Appendix B for detailed formulas for RAD-RWKV6 and RAD-RWKV7.
The protocol for converting a multilayer transformer model from one sequence mixing operation to another can be described as 3-step process:
Setup: Attention Weights Transfer. All attention related weights () from the teacher model are transferred to the student model.
Step 1: Attention Hidden State Alignment. Each student sequence mixing layer is trained to approximate the corresponding teacher model attention layer’s hidden state outputs. This can be done sequentially or in parallel.
Step 2: Knowledge Distillation. All model layers are trained together to approximate original model de-embedding output logits via Kullback-Leibler Divergence Loss.
Step 3: Fine Tuning. The new model is fine-tuned at longer contexts.
4.1 Setup: Attention Weights Transfer
When equivalent parameters are available (e.g. query = receptance, key = key, value = value, output = output) we initialize them to the teacher weights, and otherwise to a standard pretraining initialization (e.g. in RWKV6 or RWKV7 models). Some weights, such as tokenshift, are set such that they mimic the teacher model and have no immediate effect. The goal is to initially match the teacher but learn new more expressive behavior as needed throughout distillation.
4.2 Step 1: Attention Hidden State Alignment
Following notation from MLPmixer (Tolstikhin et al., 2021), we divide the layers of traditional transformers into pointwise operations (MLPs) and sequence mixing operations (Attention layers in standard transformers). In Step 1, we hold a frozen copy of the full teacher model in memory, and add a second trainable attention-replacement layer in parallel alongside each teacher attention layer, created with the replacement sequence mixing operation of choice (ex. RAD-RWKV6 or RAD-RWKV7). For loss, we employ an L2 distance (or optionally, mean squared error) objective between the student and teacher sequence-mixer hidden state outputs.
As noted earlier, this step can be done sequentially layer by layer, or in parallel with all layers at once. In practice, we train all layers at once using a single optimizer and whole model forward/backward pass.
The sequence length of this step is usually 512. We find that 100M tokens is enough to converge to a low stable loss during step 1. We run this step with a cosine annealed learning rate beginning at 1e-3 and ending at 1e-5. The final learning rate should be similar to the final learning rate the teacher saw during pretraining.
At the end of this step we remove the teacher attention layers from the model, leaving only a complete student model.
4.3 Step 2: Knowledge Distillation (Modelwise Approximation of Logits)
During this step, we load a separate copy of the full teacher model and train the student model to approximate the logits of the teacher model in a similar fashion to traditional teacher-student knowledge distillation training. (Hinton et al., 2015) We use loss equal to the Kullback-Leibler divergence of the student and teacher logits.
Empirically, we find that the model resulting from this step reproduces most of the initial model performance across a variety of benchmarks.
The sequence length of this step is usually 512. We run this step with a flat learning rate of 1e-5. This should be similar to the final learning rate the teacher saw during pretraining.
We find that 250M tokens is generally sufficient, and that training more than around 500M tokens during step 2 does not improve english language benchmarks. It’s possible that less frequently encountered tokens still improve beyond this point. This number is somewhat dependent on the quality of the attention replacement RNN used. For example, RAD-RWKV7 reaches better alignment during step 1 than RAD-RWKV6, and can therefore tolerate shorter step 2 training runs.
4.4 Step 3: Context Length Extension
During this step we train the model on longer sequences with the objective of enhancing its long-context modeling capabilities. We use regular cross-entropy loss on target labels with no teacher model. This process is done as a separate step for efficiency and memory reasons, instead of simply increasing the context length during steps 1 & 2. Steps 1 & 2 require part or all of both teacher and student models to be held in memory at the same time, while this is no longer the case in step 3. Step 3 also runs faster since it has only one model to execute, making it a better candidate for longer context lengths. It is possible to reduce the memory usage in step 2 by saving teacher logits to disk or having them generated by a separate machine, but the amount of VRAM saved would be relatively modest, since the teacher model requires neither optimizer states nor backpropagation.
We run this step with the same flat learning rate of 1e-5 as step 2.
We provide an alternative method for step 3 in situations where VRAM is at a premium. In this alternative step 3a, we freeze all weights except for decay and tokenshift (if present), and increase the learning rate to .
Name lmbda mmlu arc_c arc_e hella piqa winog Mistral-7B-v0.1-SUPRA 0.913 0.216 0.722 0.911 0.930 0.937 0.846 Mamba2-Llama3.0-8B-Instruct 0.608 0.519 0.724 0.868 0.901 0.903 0.391 MOHAWK-Phi1.5-1.3B 0.938 -0.047 0.830 0.968 0.936 0.959 0.927 DiJiang-Llama2.0-7B111Not Available: Neither the model weights nor the missing scores were published by the model authors. - 0.720 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.973 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.732 0.906 0.975 - Llamba-8B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. 0.951 0.837 0.983 1.014 0.969 0.994 0.975 LOLCatsHedgehog-Llama3.0-8B111Not Available: Neither the model weights nor the missing scores were published by the model authors. - -0.029 0.534 0.864 0.750 0.886 0.143 ARWKV-7B 0.893 0.801 0.906 0.969 0.934 0.993 0.874 Our RADLADS models: QRWKV6-7B-Instruct 0.970 0.871 1.043 0.998 0.974 1.030 1.007 QRWKV6-7B-Instruct RoPE 0.969 0.880 1.054 1.003 0.971 1.018 0.981 QRWKV7-7B-Instruct 0.982 0.924 1.026 1.004 0.979 1.013 1.041 QRWKV7-7B-Instruct-from72B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. 1.016 0.893 1.071 1.033 0.959 1.026 1.139 QRWKV6-32B-Instruct 0.986 0.909 1.066 1.037 0.964 1.005 1.235 QRWKV6-QwQ-32B 0.991 0.899 1.025 0.993 0.982 1.020 1.135 QRWKV6-72B-Instruct 1.004 0.899 1.015 1.008 0.974 0.968 1.123
Name lmbda mmlu arc_c arc_e hella piqa winog Mistral-7B-v0.1-SUPRA 0.692 0.331 0.458 0.759 0.771 0.801 0.703 teacher: Mistral-7B-v0.1 0.758 0.624 0.538 0.809 0.810 0.821 0.740 Mamba2-Llama3.0-8B-Instruct 0.439 0.452 0.480 0.741 0.708 0.758 0.586 teacher: Llama3.0-8B-instruct 0.722 0.639 0.567 0.816 0.758 0.786 0.719 MOHAWK-Phi1.5-1.3B 0.501 0.242 0.441 0.740 0.602 0.755 0.717 teacher: Phi1.5-1.3B-base 0.534 0.418 0.480 0.756 0.626 0.766 0.734 DiJiang-Llama2.0-7B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. - 0.407 0.427 0.626 0.694 0.775 - teacher: Llama2.0-7B - 0.468 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.463 c0.764 0.740 0.782 - Llamba-8Ba 0.694 0.610 0.546 0.825 0.776 0.809 0.733 1st teacher: Llama3.1-8B-Instruct 0.730 0.680 0.551 0.817 0.793 0.811 0.739 2nd teacher: Llama3.1-70B-Instruct 0.757 0.823 0.624 0.867 0.847 0.837 0.787 LoLCATsHedgehog-Llama3.0-8B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. - 0.238 0.401 0.726 0.656 0.765 0.533 teacher: Llama3.0-8B 0.666 0.533 0.801 0.791 0.799 0.731 0.381 ARWKV-7B 0.622 0.624 0.522 0.797 0.768 0.792 0.687 QRWKV6-7B-Instruct 0.675 0.657 0.563 0.814 0.790 0.803 0.711 QRWKV6-7B-Instruct-RoPE 0.674 0.661 0.567 0.817 0.789 0.799 0.706 QRWKV7-7B-Instruct-RoPE 0.684 0.682 0.558 0.817 0.793 0.798 0.718 teacher: Qwen2.5-7B-Instruct 0.696 0.717 0.550 0.815 0.805 0.794 0.714 QRWKV7-7B-Instruct-from72B111Not Available: Neither the model weights nor the missing scores were published by the model authors. 0.707 0.667 0.572 0.833 0.782 0.801 0.739 teacher: Qwen2.5-72B-Instruct 0.751 0.834 0.632 0.859 0.874 0.836 0.763 QRWKV6-32B-Instruct 0.742 0.766 0.609 0.843 0.830 0.812 0.782 teacher: Qwen2.5-32B-Instruct 0.752 0.818 0.587 0.822 0.852 0.810 0.729 QRWKV6-QwQ-32B 0.662 0.743 0.564 0.806 0.830 0.804 0.732 teacher: Qwen2.5-QwQ-32B 0.668 0.799 0.556 0.810 0.841 0.798 0.705 QRWKV6-72B-Instruct 0.754 0.775 0.638 0.865 0.857 0.825 0.796 teacher: Qwen2.5-72B-Instruct 0.751 0.834 0.632 0.859 0.874 0.836 0.763
5 Language Modeling Performance
In Table 3 we compare pure RNN models output from various conversion methods across a set of standard benchmarks. These converted models were trained from variety of different teacher models and vary in terms of parameter count. Therefore, we compare them via relative score: the ratio of benchmark accuracy scores between each student and teacher model, first subtracting the random chance accuracy for each benchmark. Negative ratios indicate that a model performed worse than random chance. Note that on nearly every benchmark, the RADLADS ”QRWKV” models obtain higher score ratios than models converted via any other method. (”QRWKV” is our shorthand for ”Qwen with RWKV attention”) In Table 3 QRWKV6-72B-Instruct demonstrates a new state-of-the-art in downstream performance for a pure RNN language model.
Next, in Tables 5 and 5 we show models converted to hybrid-attention variants, each containing some amount of softmax attention. Even when compared to these hybrid models that still employ softmax attention, the pure RNN QRWKV models show the highest MMLU relative scores, and achieve similar or better relative scores on other benchmarks.
Name Attention mmlu arc_c arc_e hella piqa winog LoLCATs-Mistralv0.1-7B SWA 100% 0.706 1.038 1.014 0.995 0.981 1.000 LoLCATs-Llama3.1-8B SWA 100% 0.727 1.037 1.015 1.001 1.069 0.839 LoLCATs-Llama3.1-70B SWA 100% 0.794 0.997 0.963 0.994 0.970 0.801 Mamba2-Llama3.0-8B-Instruct 12.5% 0.663 0.797 0.898 0.948 0.937 0.611 Mamba2-Llama3.0-8B-Instruct 25% 0.738 0.867 0.907 1.037 1.004 0.676 Mamba2-Llama3.0-8B-Instruct 50% 0.789 1.047 0.951 1.072 1.099 0.982
Name Attention mmlu arc_c arc_e hella piqa winog LoLCATs-Mistralv0.1-7B SWA 100% 0.514 0.549 0.817 0.807 0.815 0.740 teacher: Mistralv0.1-7B 100% 0.624 0.538 0.809 0.810 0.821 0.740 LoLCATs-Llama3.1-8B SWA 100% 0.549 0.544 0.824 0.791 0.810 0.697 teacher: Llama-3.1-8B 100% 0.661 0.534 0.815 0.790 0.790 0.735 LoLCATs-Llama3.1-70B SWA 100% 0.677 0.605 0.850 0.846 0.821 0.737 teacher: Llama-3.1-70B 100% 0.788 0.606 0.873 0.850 0.831 0.796 Mamba2-Llama3.0-8B-Instruct 12.5% 0.508 0.503 0.758 0.731 0.768 0.634 Mamba2-Llama3.0-8B-Instruct 25% 0.537 0.525 0.764 0.777 0.787 0.648 Mamba2-Llama3.0-8B-Instruct 50% 0.557 0.582 0.788 0.795 0.815 0.715 teacher: Llama3.0-8B-Instruct 100% 0.639 0.567 0.816 0.758 0.786 0.719
6 Ablation Studies
We performed several ablation studies, adding or removing individual mechanisms to the RAD-RWKV6 architecture at the 7B scale. For the ”use groupnorm” ablation we added the use of GroupNorm and removed the state balancing mechanism originally from RWKV6-C2. Results are shown in Table 6.
Arch ablation lambada MMLU arc_c arc_e hella piqa winog RAD-RWKV6 none 0.6748 0.6572 0.5631 0.8136 0.7901 0.8025 0.7111 RAD-RWKV6 use rope 0.6740 0.6610 0.5666 0.8165 0.7886 0.7992 0.7056 RAD-RWKV6 no tokenshift 0.6707 0.6584 0.5546 0.8043 0.7874 0.8036 0.6875 RAD-RWKV6 no gate 0.6590 0.6417 0.5444 0.8077 0.7842 0.7971 0.6930 RAD-RWKV6 use groupnorm 0.6559 0.6340 0.5648 0.8186 0.7865 0.7905 0.7032
7 Training Details and Hyper-parameters
We use the DCLM (Li et al., 2024b) dataset for all three steps in the RADLADS attention conversion process.
Step Tokens LR LR decay seqlen batch size adam betas, eps 1 100M cosine 512 32 0.9, 0.95, 2 250M-700M none 512 96 0.9, 0.95, 3 100M none 16384 96 0.9, 0.95,
Model Step 1 Step 2 Step 3 Size GPU Tokens Opt Hours Tokens Opt Hours Tokens Opt Hours 7B 8x Mi300X 100M DS1 0.75 500M DS1 5.5 100M DS1 1 32B 8x Mi300X 100M DS1 2.5 500M FSDP 27.0 100M FSDP 3 72B 8x Mi300X 100M FSDP 7.50 500M FSDP 54.0 100M FSDP a6
7.1 General
We manually choose between DeepSpeed (Li et al., 2024a) ZeRO stages 1 or 2 and FSDP (Zhao et al., 2023), depending on VRAM availability. Typically we use DeepSpeed Zero stage 1 for our step 1 whenever possible, as it results in faster training and step 1 requires less VRAM than step 2. We find that FSDP is more VRAM efficient than DeepSpeed ZeRO Stage 3 for larger models, and we generally use it for steps 2 and 3.
8 What Did Not Work
In the spirit of scientific transparency and to facilitate reproducibility and further experimentation in the field, we share some negative results from our experiments:
Initial Attention Score Alignment
We originally included an initial ”step 0”, which was similar to step 1 but used a loss based on attention scores differences. For linear RNNs of the form
attention scores can be represented as cumulative compositions of transforms:
This generalizes Mamba2 (Dao & Gu, 2024), RWKV5, RWKV6 (Peng et al., 2024), RWKV7 (Peng et al., 2025), DeltaNet (Schlag et al., 2021), and Gated DeltaNet (Yang et al., 2025), as well as others. See Table 1 in Peng et al. (2025) for a detailed list of such RNNs.
We hypothesized that matching attention scores up-front would accelerate training compared to Step 1. While final downstream performance was similar, we observed neither benefits in total training time nor lower final loss. On the other hand, extended training of step 0 incurred final higher loss.
Skipping step 1
Starting directly with step 2 distillation and skipping step 1 entirely resulted in a much lower performance model. The same level of convergence simply did not occur, with loss plateauing at a higher minimum, even with longer training.
De-novo initialization of attention weights
Initializing sequence mixer QKVO weights as if they were untrained, instead of copying them from the teacher model, resulted in consistently worse yet surprisingly reasonable performance. This is compatible with the previous hypothesis stating that factual knowledge is stored mainly in MLP model weights.
Freezing model weights
Although it might seem intuitive that the MLPs and embeddings could be frozen during step 2, we find that this results in significantly reduced model performance. We theorize that the internal embedding and hidden state representations need to adapt somewhat to the new RNN sequence mixer’s use of channels. Through continued training, the MLPs may learn to route this information differently to avoid conflicts with pre-existing information flow from the teacher model.
Larger batchsizes
The number of optimizer steps appears to be of key importance during conversion. Consequently, increasing batchsize did not help the model converge faster. Additionally, we are careful to keep the learning rate low during steps 2 and 3 in order to avoid excess disturbance to the MLPs. Therefore, adjusting the learning rate higher to compensate for larger batchsizes would be undesirable.
LoRA training
We tried using LoRA to perform PEFT, but rank reduction was generally quite detrimental to performance. The one place we found it could work without causing problems was on the embeddings.
Switching datasets
We tried using various custom datasets during the context-length extension step. These seemed to create a confused model that would use more and more adjectives as generation progressed. We leave exploration of the ideal dataset for this step to future work.
9 Conclusions
We have shown that RADLADS provides a cost-effective way to convert transformers employing quadratic softmax attention into inexpensive RNN models that feature linear time complexity and constant memory usage. This has potential to save energy, reduce research costs, and enable the rapid testing and deployment of new varieties of RNN architectures. We have already used it to produce and distribute new state-of-the-art open weight linear attention models.
However, RADLADS does have known limitations. Currently, each new architecture design requires meticulous testing to improve its compatibility with the RADLADS protocol. As one example, RAD-RWKV7 appears to exhibit reduced training stability at larger (32B+) parameter and layer counts, and we are actively working to design architectural and training methodology changes to remedy this. New varieties of model interactions will test these conversions in ways that we cannot predict. For example, the details of how the conversion process impacts reasoning models like our converted QwQ model is still unknown and requires further testing.
We have many ideas for future work to enhance and extend the RADLADS process. Converting to and from different architectures and testing of more varieties of datasets against each of these would help discover the principles for optimal conversion dataset design. We also believe that parts of our existing RAD-RWKV7 architecture can still be improved; balancing of state to create implicit normalization and then removal of groupnorm could further increase downstream performance.
Acknowledgments
Eric Alcaide acknowledges support for this work from the Swiss State Secretariat for Education, Research and lnnovation (SERI) under contract number 23.00421.
Appendix A Author Contributions
Appendix B Architecture Detail
Notation
In this section, we use to denote the model dimension. Bold capital letters represent trainable matrices, and vectors without a subscript are trainable parameters. The first subscript denotes sequence position and second subscript denotes layer index, where necessary. We use the convention that all vectors are row vectors unless explicitly transposed, so all matrices operate on the right side, therefore is an outer product and is an inner one. We use the square subscript to denote a placeholder for variable names and use the sign for cumulative matrix multiplication.
B.1 RAD-RWKV6 Time Mixing
RAD-RWKV6 is a customized variation of RWKV6-C2 (”Finch-C2”) (Goldstein et al., 2024), featuring a Gated Linear Attention (Yang et al., 2024) kernel (so no bonus), and a sigmoid gate. The use of the state balancing technique from RWKV6-C2 enables us to remove state normalization, improving downstream performance. Removal of bonus did not harm downstream performance. We do not typically use RoPE in this formulation.
The data-dependent linear interpolation (ddlerp) between and used in Finch Token Shift is defined as:
where and each introduce a trainable vector and and each introduce a pair of low rank trainable matrices, where is a chosen size for each such pair.
| receptance | (3) | ||||
| value | (4) | ||||
| gate | (5) | ||||
| decay precursor | (6) | ||||
| decay | (7) | ||||
| key precursor | (8) | ||||
| key | (9) |
Time mixing is performed by the attention calculation:
The attention calculation can alternatively be written in a recurrent manner:
Finally, the heads are recombined via reshaping so that , gated, and transformed into the output as follows:
B.2 RAD-RWKV7 Time Mixing
RAD-RWKV7 is a customized variation of RWKV-7 (Peng et al., 2025), with tokenshift removed, RoPE applied, and no bonus. The removal of tokenshift speeds up training and inference, and bonus had no beneficial impact on downstream performance.
| value residual gate | (21) | ||||
| value precursor | (22) | ||||
| decay precursor | (24) | ||||
After weight preparation, we reshape , splitting them to heads, with each head sized . We always assume that is a factor of and heads are equally split. All operations in this section are shown per-head.
Before mixing in the time dimension, is normalized per head:
| normalized removal key | (28) |
Time mixing is performed by the attention calculation:
The attention calculation can alternatively be written in a recurrent manner:
| attention result | (32) |
Finally, the heads are recombined via reshaping so that , gated, and transformed into the output as follows: